**Please answer if you really know, really wanna learn from the answers, THANKS!**

A cube of side length L = 0.11 m occupies a volume of space in which the electric field along the positive z-axis is 4.9 x 103 z2 N/C. Along the x- and y-axes the field components are uniform, equal to 1,000 N/C for both. What is the total (net) charge inside the cube? Give your answer in pC (1 pC = 10-12 C) but enter only the numerical part.

N-
Answer #1

p= Q = $E-da | E= 1000 +1000+4 1 310k HEDA - SS6000 dydz + 1000 dudz +4.9x10dxdy) Q - (1000 +1000+4.9x10°)x 0112 Q = to (83.